OpenStudy (anonymous):

Find the coordinates of the midpoint between point A(-5, -9) and the origin. A. (5/2, 9/2) B. (-5, -9) C. (-5/2, -9/2 D. (5, 9) I don't understand this at all :S

OpenStudy (shane_b):

Use the midpoint formula:\[(\frac{x_1+x_2}{2},\frac{y_1+y_2}{2})\]Where x1 and y1 = 0, x2 =-5 and y2=-9.

OpenStudy (shane_b):

You're basically dividing the distances between each point on each axis in half...that's the easy way to remember the formula.